OpenStudy (anonymous):

An 8-ounce serving of soup contains 19 grams of carbohydrates. To the nearest gram, how many grams of carbohydrates does a 5-ounce serving of soup contain?

OpenStudy (anonymous):

A. 10 grams B. 11 grams C. 12 grams D. 13 grams Which one of these answers is correct?

OpenStudy (bibby):

Set up a proportion.\[\huge \frac{ 8oz }{ 19g } = \frac{ ? }{ ? }\] what goes on the right hand side?

OpenStudy (anonymous):

I know i have to divide but i dont what to divide it with

OpenStudy (bibby):

In proportions, you want to keep the units consistent across the fractions. That means ouncez on top and gramz on bottom

OpenStudy (anonymous):

So is it 1/2?

OpenStudy (bibby):

I'm talking about the question marks \[\huge \frac{ 8oz }{ 19g } = \frac{ ? }{ ? }\]you can't divide anything before filling out the equation

OpenStudy (anonymous):

Oh, 19 and 5 right?

OpenStudy (bibby):

Reread the question. 8-ounce soup contains 19 grams of carbs. how many does a 5-ounce serving of soup contain? That means we don't know one of the numbers.

OpenStudy (anonymous):

oh okay

OpenStudy (bibby):

How much more direct can I make this? 8-ounce soup = 19 grams of carbs. 5-ounce soup = ? carbs